A Bibliography of Publications of Makan Pourzandi

Makan Pourzandi Lab. LIP-IMAG ENS Lyon, 46 Allee d’Italie 69364 Lyon Cedex 07 FRANCE Tel: (+33) 72 72 85 03 FAX: (+33) 72 72 80 80 E-mail: mpourzan@lip.ens-lyon.fr (Internet) 25 November 2011 Version 1.03

Abstract

This bibliography records publications of Makan Pourzandi.

Title word cross-reference

860 [PT93a]. algorithms [DP94]. comparison [DP94]. Diffusion [PT93b]. eigenvalue [PT94]. Elimination [PT93a]. Equation [PT93b]. Gaussian [PT93a]. iPSC [PT93a]. iPSC/860 [PT93a]. IS-860 [PT93b].

Jacobi [PT94]. Jacobi-like [PT94].

like [PT94]. matrix [DP94]. Overlapping [PT93a]. parallel [DP94]. performance [PT94]. Preliminary [PT93b]. product [DP94]. Resolution [PT93b]. solution [PT94]. Stochastic [PT93b]. Study [PT93b]. System [PT93b]. three [DP94]. VolVox [PT93b].

References

Desprez:1994:CTP

[DP94] F. Desprez and M. Pourzandi. A comparison of three parallel matrix product algorithms. In proceedings of the third International Conference on Numerical Methods and Applications. Bulgarian Academy of Sciences, Bulgaria, August 1994.

Pourzandi:1993:OGE

[PT93a] M. Pourzandi and B. Tourancheau.

Overlapping in Gaussian elimination on iPSC/860. In M. A. Yaghoubi, editor, Proceeding of International Congress on Computational Methods in Engineering, volume 4, pages 183– 193. University of Shiraz, Iran, May 1993.

Pourzandi:1993:SRD

[PT93b] M. Pourzandi and B. Tourancheau.

Stochastic Resolution of a Diffusion Equation System on VolVox IS-860, preliminary study. Technical Report 93-31, Laboratoire de l’Informatique du Parallelisme - Ecole Normale Supérieure de Lyon, October 1993.

Pourzandi:1994:PPS

[PT94] M. Pourzandi and B. Tourancheau.

A parallel performance study of Jacobi-like eigenvalue solution. Technical Report 94-16, Laboratoire d’Informatique du Parallélisme, March 1994. Also disponible as technical report at Computer Science Department of University of Tennessee.